OpenStudy (anonymous):

Anyone know Spearman's Rank correclation coefficient? I need help with a problem...

OpenStudy (anonymous):

Refer this it might be helpful http://en.wikipedia.org/wiki/Spearman%27s_rank_correlation_coefficient

OpenStudy (anonymous):

I know that it's \[\rho = 1- {\frac {6 \sum d_i^2}{n(n^2 - 1)}}. \] But I'm solving an equation... and the answer at the back of the book isn't the same as my answer, as long as I've been trying it, and I don't know what I'm doing wrong... so I was wondering if anyone was familiar with it to see if they got the same answer as that of the back of the book?

OpenStudy (phi):

What are the details? Do you have x and y values?

OpenStudy (anonymous):

Yes, I can give them if you know how to use it...

OpenStudy (anonymous):

rs=1-6summitionDsquare/ncube-n

OpenStudy (anonymous):

Ph k sry then give the details and I will try my best

OpenStudy (anonymous):

Magazine 1: 13 20 22 18 19 11 10 15 Magazine 2: 17 19 23 16 20 10 11 18

OpenStudy (phi):

First sort the 2 sequences, and number them 1..N

OpenStudy (anonymous):

dind uot difference

OpenStudy (anonymous):

and then do the square of the difference

OpenStudy (anonymous):

Yes, That's what I thought. What do you mean by sorting? Aren't they already sorted?

OpenStudy (anonymous):

here n=8

OpenStudy (anonymous):

assign them rank

OpenStudy (anonymous):

how do you assign them rank?

OpenStudy (phi):

m1 10, 11, 13, 15, 18, 19, 20, 22 1 2 3 4 5 6 7 8 m2 10, 11, 16, 17, 18, 19, 20, 23 1 2 3 4 5 6 7 8 Magazine 1: 13 20 22 18 19 11 10 15 3 7 8 4 5 2 1 3 Magazine 2: 17 19 23 16 20 10 11 18 3 5 8 3 6 1 2 4

OpenStudy (phi):

Sorted mag 1's numbers, and assigned it 1,2,3 ... from small to large. Used these to label the original sequence.

OpenStudy (phi):

now subtract the ranks to get the d's

OpenStudy (phi):

Magazine 1: 13 20 22 18 19 11 10 15 3 7 8 4 5 2 1 3 Magazine 2: 17 19 23 16 20 10 11 18 3 5 8 3 6 1 2 4 d's 0 2 0 1 -1 1 -1 -1 d sq 0 4 0 1 1 1 1 1

OpenStudy (anonymous):

How did you get that? I mean the d's

OpenStudy (phi):

First, I labeled the 13 of mag 1 with its rank (3, which is its order if you sorted mag 1's numbers). I labeled mag 2's 17 with a 3 (its rank) d= 3-3= 0 ditto for the rest...

OpenStudy (anonymous):

I thought 17 was 4? You don't have a 7 in your ranking in magazine 2..?

OpenStudy (phi):

Yes, you are right... Feel free to fix up my mistakes.

OpenStudy (anonymous):

Ok... If I get the answer now, I'll tell you...

OpenStudy (anonymous):

I got it :) Thanks!

OpenStudy (phi):

If I didn't make any mistakes Magazine 1: 13 20 22 18 19 11 10 15 3 7 8 5 6 2 1 4 Magazine 2: 17 19 23 16 20 10 11 18 4 6 8 3 7 1 2 5 d's -1 -1 0 2 -1 1 -1 -1 d sq 1 1 0 4 1 1 1 1

OpenStudy (anonymous):

Yes that's right!

OpenStudy (phi):

So the coeff is about 0.88